OpenStudy (batman31):

I am doing this experiment, and have reached the point where you apply snell's law, and I don't get it. Link is in the comments.

OpenStudy (anonymous):

LINK

OpenStudy (batman31):

number 9.

OpenStudy (batman31):

how does this equation work? walk me through it

OpenStudy (anonymous):

you need the numbers that go in the equation ...

OpenStudy (batman31):

27 and 17

OpenStudy (anonymous):

which is sin and which is V

OpenStudy (batman31):

neither, from this experiment I do not know any values for 4.

OpenStudy (batman31):

the WIP equation is \[(\sin 17)/(\sin 27)=v1/v2\]

OpenStudy (anonymous):

@Nnesha @iGreen

OpenStudy (batman31):

I just watched the khan academy vid on snell's law, and it didn't clear much up. @directrix

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

yeah i dont understand

OpenStudy (batman31):

@Michele_Laino

OpenStudy (michele_laino):

here is a little explanation:  the statement of the Snell's law is: \[\Large {n_1}\sin {\theta _1} = {n_2}\sin {\theta _2}\]

OpenStudy (michele_laino):

since the light speed v in amedium with refractive index n, is: \[\Large v = \frac{c}{n}\] then we can divide both sides of the Snell's law by c, so we get: \[\Large \frac{{{n_1}}}{c}\sin {\theta _1} = \frac{{{n_2}}}{c}\sin {\theta _2}\]

OpenStudy (michele_laino):

now, according to the previous formula, we have: \[\Large {v_1} = \frac{c}{{{n_1}}},\quad {v_2} = \frac{c}{{{n_2}}}\]

OpenStudy (michele_laino):

then the statement of the Snell's law can be rewritten as follows: \[\Large \frac{{\sin {\theta _1}}}{{{v_1}}} = \frac{{\sin {\theta _2}}}{{{v_2}}}\] and finally: \[\Large \frac{{\sin {\theta _1}}}{{\sin {\theta _2}}} = \frac{{{v_1}}}{{{v_2}}}\]

OpenStudy (batman31):

@Michele_Laino are you still here? I still don't get it. I was eating

OpenStudy (batman31):

@paki @texaschic101

OpenStudy (michele_laino):

why?

OpenStudy (michele_laino):

please, read my step-by-step explanation